A university's leadership council is made up of 14

 members: 5

 are students, 4

 are teachers, and 5

 are deans. A newspaper reporter is planning to select two of the council members for interviews. The first member will be chosen at random, and then the second member will be chosen at random from the remaining members. What is the probability that two teachers will be selected?


1
Expert's answer
2021-05-11T09:22:20-0400

"P=\frac{4}{14}*\frac{3}{13}=\frac{6}{91}=0.0659."